# KCL-PH-TH/2026-03

### v2

@@ II The Effective Action of Heterotic String Theory
$\begin{split}S_{\text{GCS}}&=\int\Bigg\{\frac{1}{16\kappa^{2}}\left(R_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}-4R_{\mu\nu}R^{\mu\nu}+R^{2}\right)\\ &\phantom{\int}+\left(-\frac{1}{8}+\frac{1}{24\sqrt{3}}\right)\mathcal{H}_{\mu\nu\kappa}\mathcal{H}_{\lambda\rho}{}^{\kappa}R^{\mu\nu\lambda\rho}+\left(\frac{1}{4}-\frac{1}{24\sqrt{3}}\right)\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}-\frac{1}{24}\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}R\\ &\phantom{\int}+\kappa^{2}\left[\frac{1}{24}\left(1-\frac{2}{9\sqrt{3}}\right)\right]\mathcal{H}^{\mu\nu\lambda}\mathcal{H}_{\mu}{}^{\rho\sigma}\mathcal{H}_{\nu\rho}{}^{\kappa}\mathcal{H}_{\lambda\sigma\kappa}+\kappa^{2}\left(-\frac{1}{24}+\frac{\sqrt{3}}{648}\right)\mathcal{H}^{\mu\lambda\rho}\mathcal{H}_{\nu\lambda\rho}\mathcal{H}^{\nu\kappa\sigma}\mathcal{H}_{\mu\kappa\sigma}\\ &\phantom{\int}+\kappa^{2}\left(\frac{1}{144}\right)(\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda})^{2}\Bigg\}\sqrt{-g}\,d^{D}x.\end{split}$ (39)
Comparison with (16) shows that no choice of $B_{3},B_{4}$ can yield (39) (a particularly easy way to see this is to try to match both $\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}$ and $\mathcal{H}^{\mu\lambda\rho}\mathcal{H}_{\nu\lambda\rho}\mathcal{H}^{\nu\kappa\sigma}\mathcal{H}_{\mu\kappa\sigma}$. This, then, leads to contradicting solutions for the value of $B_{3}$). In other words, in $D>4$ dimensions we have to choose between a unitary action and interpreting the Kalb-Ramond field strength as torsion [30, 31]. However, this is not necessarily the case in $D=4$. There, due to over-antisymmetrization, the following dimensionally dependent identities hold:
@@
One way to confirm these identities is to replace $\mathcal{H}$ with its dual in $D=4$, i.e.
@@
We may use these identities to simplify the actions (16) and (32) (post-compactification to (3+1)-dimensions) and obtain [^7] It should be noted that, since the field redefinitions are performed in the $D=10$-dimensional low-energy heterotic string effective action, we should set $D=10$ in equation (16). Then, we may proceed with the compactification to $D=4$ dimensions, and only afterwards we apply the dimensionally dependent identities to the (3+1)-dimensional action. Note that the processes of field redefinition and compactification do not “commute”, in the sense that if we compactify first and then do the field redefinitions in $D=4$, we obtain an action with different coefficients. For more details, see Appendix (B) :
@@
and [^8] Interestingly, the invariant $\bar{G}^{+}_{6}$ vanishes in $4D$ for a totally antisymmetric torsion. The invariant $\bar{G}^{+}_{4}$, being the Gauss-Bonnet combination with torsion (Euler density in Einstein-Cartan spacetimes), reduces to the ordinary Gauss-Bonnet combination in the absence of torsion (Euler density in Einstein (General Relativity) spacetimes).
$\begin{split}S_{\text{GCS}}^{D=4}&=\int\bigg\{\frac{1}{\kappa^{2}}(A_{4}+A_{5})(R_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}-4R_{\mu\nu}R^{\mu\nu}+R^{2})+\frac{8}{3}A_{5}\left(\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}-\frac{1}{3}\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}R\right)\bigg\}\sqrt{-g}\,d^{4}x.\end{split}$ (45)
@@
$\begin{split}S&=\int\left[\frac{1}{2\kappa^{2}}\bar{R}+\frac{\alpha^{\prime}}{\kappa^{2}}\left(A_{4}\bar{G}^{+}_{4}+A_{5}\bar{G}^{+}_{5}\right)\right]\sqrt{-g}\,d^{4}x\\ &=\int\left[\frac{1}{2\kappa^{2}}\bar{R}+\frac{\alpha^{\prime}}{\kappa^{2}}\left[A_{4}\left(\bar{R}_{\mu\nu\lambda\rho}\bar{R}^{\lambda\rho\mu\nu}-4\bar{R}_{\mu\nu}\bar{R}^{\nu\mu}+\bar{R}^{2}\right)+A_{5}\left(\bar{R}_{\mu\nu\lambda\rho}\bar{R}^{\mu\nu\lambda\rho}-4\bar{R}_{\mu\nu}\bar{R}^{\mu\nu}+\bar{R}^{2}\right)\right]\right]\sqrt{-g}\,d^{4}x\\ &=\int\bigg[\frac{1}{2\kappa^{2}}\bar{R}+\frac{\alpha^{\prime}}{\kappa^{2}}\bigg[\frac{1}{16}\bar{R}^{2}+\frac{3}{224}\left(\frac{17}{3}-\frac{22}{9\sqrt{3}}\right)\left(\bar{R}_{\mu\nu\lambda\rho}\bar{R}^{\lambda\rho\mu\nu}-4\bar{R}_{\mu\nu}\bar{R}^{\nu\mu}\right)\\ &\qquad\qquad\qquad\qquad\qquad\quad-\frac{3}{224}\left(1-\frac{22}{9\sqrt{3}}\right)\left(\bar{R}_{\mu\nu\lambda\rho}\bar{R}^{\mu\nu\lambda\rho}-4\bar{R}_{\mu\nu}\bar{R}^{\mu\nu}\right)\bigg]\bigg]\sqrt{-g}\,d^{4}x\end{split}$ (51)
In $D=4$, the Gauss-Bonnet combination is a total derivative, which we may ignore. Thus, the full effective action reads:
$S=\int\left[\frac{1}{2\kappa^{2}}R-\frac{1}{6}\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}+\alpha^{\prime}\left(\frac{11}{126\sqrt{3}}-\frac{1}{28}\right)\left(\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}-\frac{1}{3}\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}R\right)\right]\sqrt{-g}\,d^{4}x.$ (52)
Since $\bar{G}^{+}_{4}$ is the Gauss-Bonnet combination in an Einstein-Cartan manifold, it is easy to see, from (45) and (51), that the effective string action is characterized by a single quadratic curvature invariant, namely $\bar{G}^{+}_{5}$, which provides the non-vanishing $\mathcal{H}\mathcal{H}R$ terms. In the next section we shall discuss the emergence of dynamical axion fields from the totally antisymmetric torsion interpretation of the Green-Schwarz-modified Kalb-Ramond field strength (4), following [23] (or [2] in the StRVM context).
@@ IV Axion-like fields from the $\mathcal{H}$-Torsion
To this end, we first rewrite the $D=4$ effective action (52) in a more conventient form, with generic coefficient of the higher-order corrections, as:
@@
$A=\alpha^{\prime}\left(\frac{11}{128\sqrt{3}}-\frac{1}{28}\right)\simeq 0.015\,\alpha^{\prime}\,$, (54)
in the heterotic string case.
In $D=4$, the Bianchi identity (7) can be contracted with the Levi-Civita tensor $\eta_{\mu\nu\lambda\rho}$, so that it can be expressed, in tensor notation, as the scalar identity
@@
The Levi-Civita tensor in $D=4$ is defined as [24]
@@
where $\epsilon$ denotes the Levi-Civita symbol (tensor density) with the convention $\epsilon_{0123}=+1$, and $\epsilon^{0123}=-1,$ etc. Furthermore, the symbol (~)($\widetilde{\cdots}$) over the curvature and gauge field-strength tensors denotes the corresponding duals, defined as
@@
We can now impose the Bianchi identity (55) as a constraint in the path integral of the action (53), following [23, 2]. We do this by inserting the identity as a delta functional in the path integral:
@@
We can express this delta functional as a path integral of a pseudoscalar Lagrange multiplier field b(x)b(x): [^9] The partition function (58) is parity ($P$) even, while the CS anomaly $\mathcal{G}(\boldsymbol{\omega},\boldsymbol{A})$ is parity odd. Hence, the field $b(x)$ must be a pseudoscalar so that the right-hand-side of (59) below is parity even, thus matching the even-parity nature of its left-hand side.
@@
where we have performed a partial integration to move the partial derivative to $b$ and discarded the total derivative term. Thus our effective action becomes
@@
This action is still quadratic in $\mathcal{H}$ and, therefore, can be integrated out analytically. To this end, we first write it in the form:
@@
and we can proceed to integrate out $\mathcal{H}$. We perform a Wick rotation to a Euclidean version of the path integral:
@@
and so $\det K=C\det Q$, where $C$ is a (formally infinite) field-independent “constant”, with no physical significance, which can be absorbed in the path-integral normalization constant $\mathcal{N}$, and thus canceled when one computes normalized correlation functions, which are physically relevant.
@@
such that the overall factor of $\frac{1}{3}$ gets absorbed into the path-integral normalization constant. Then, we can use the identity
@@
where the trace tr is over the internal (tensor) indices). In our case, the function $f$ is the natural logarithm and thus
@@
Thus, we have calculated the (divergent) one-loop correction at order $O(A)\sim O(a^{\prime})$ to be
@@
We can then Wick rotate back to Minkowski space and write the full path integral as:
@@
Finally, we shall determine the inverse of $K^{\mu\nu\lambda}{}_{\alpha\beta\gamma}$. Schematically, we have that,
@@
Since $A\propto a^{\prime}$, we only keep the $\propto A$ term in the expansion. Therefore, we have to this order in $\alpha^{\prime}$:
@@
Before proceeding with calculating $JK^{-1}J$, and thus the resulting effective action in terms of metric and axion $b$ fields, we should now give, for completeness, the modification of (1) due to the higher-order corrections of the effective action (53), proportional to the coefficient $A$ ((54)). This can be achieved by studying the saddle points of the corresponding path integral over the $\mathcal{H}$-field for the action (61), (62). The latter correspond to the classical equations of motion of the action with respect to the field strength $\mathcal{H}_{\mu\nu\lambda}$, which read:
@@
where $J_{\alpha\beta\gamma}$ is given in (62), and the inverse $(K^{-1})_{\mu\nu\lambda}^{\quad\,\,\,\,\alpha\beta\gamma}$ in (86), (84), to linear order in $A\propto\alpha^{\prime}$ (cf. (54)). After some straightforward algebraic manipulations, we finally arrive at:
@@
Then, we can use the Schouten identity, namely the fact that overantisymmetrized objects vanish:
@@
which is consistent with the generic result that a three form in (3+1)-dimensional spacetimes is dual to a vector $\mathcal{V}^{\mu}$, in the sense of the relation
@@
Indeed, from (IV), (91), we can calculate the dual vector as
@@
For $A=0$ this yields (1), while for $A\neq 0$ we observe that one obtains modifications of the duality relation which amount to a direction dependent term, proportional to components of the Ricci tensor contracted with $\partial^{\mu}b$. In the context of our StRVM, we are interested in (3+1)-dimensional Einstein spaces, for which $R_{\mu\nu}=\frac{R}{4}\,g_{\mu\nu}$, for which (90) reduces to
@@
which amounts to a modification of (1) by a global rescaling factor depending on the (torsion-free) scalar curvature, of the maximally-symmetric Einstein spacetime. As we shall discuss later on in the article, the quantity $|A\,R|\ll 1$ during the RVM inflationary scenario, so any singular behaviour is avoided in this case.
@@
Therefore, we have finally:
@@
The resulting effective action is therefore given by:
@@
On rescaling $b\rightarrow\frac{1}{3\sqrt{2}}b$ we finally obtain an effective action with canonical kinetic terms for the KR axion field: [^10] We follow here the conventions and normalizations of [23], which differ from those of [2] by a rescaling of $\alpha^{\prime}$ by a factor $\sqrt{3}$.
@@
On replacing $\delta^{(4)}(0)\sim\Lambda^{4}$, where $\Lambda$ is an UV cutoff energy scale, we observe that the coefficient of the Einstein-Hilbert (scalar-curvature) term in the effective action (99), is positive, and so unitarity is guaranteed.
In string effective field theories, the UV cutoff is the string mass scale,
@@
whose value is bounded from above by $M_{\rm Pl}$ [19]. Using (100), the final form of the string-inspired effective action (99), containing terms of fourth order in derivatives, and respecting unitarity and torsion interpretation of $H_{\mu\nu\rho}$, reads:
@@
Therefore, we observe that the effects of keeping the complete four-derivative $\mathcal{O}(\alpha^{\prime})$ terms in the effective action of the bosonic massless string multiplet, under the assumption of constant dilaton, are:
@@
  • Introduction of a non-minimal derivative coupling of the KR axion b with the Ricci tensor, which maintains the shift-symmetry of the axion, as it should, given that such terms originate from a path integration of the $H$-field strength, which is classically related to $b$ via (1).
It is worth noting that, when expressed in terms of the dimensionless $\widetilde{b}(x)\equiv\kappa\,b(x)$ axion field, the coefficients of the four-space-time derivative terms, that is, the CS anomalous terms and the non-minimal derivative coupling term, are both proportional to order $\alpha^{\prime}/\kappa^{2}\lesssim 1$. However, the relative magnitude of these terms can only be determined once we know the corresponding contributions at specific cosmological eras, which we shall do in the next section. Since we’re working in an EFT/perturbative approach, the ghosts introduced by both the CS and the non-minimal derivative coupling term can be ignored as long as we stay below the UV cutoff [7, 8]. We now proceed to calculate the equations of motion. It will be convenient in what follows to use the following generic form of our effective action (101):
@@
For completeness we note that, on using the Ricci-tensor identity

### v3

@@ II The Effective Action of Heterotic String Theory
$S_{\text{GCS}}$ $=\int\Bigg\{\frac{1}{16\kappa^{2}}\left(R_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}-4R_{\mu\nu}R^{\mu\nu}+R^{2}\right)$
$\phantom{\int}+\left(-\frac{1}{8}+\frac{1}{24\sqrt{3}}\right)\mathcal{H}_{\mu\nu\kappa}\mathcal{H}_{\lambda\rho}{}^{\kappa}R^{\mu\nu\lambda\rho}+\left(\frac{1}{4}-\frac{1}{24\sqrt{3}}\right)\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}-\frac{1}{24}\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}R$
$\phantom{\int}+\kappa^{2}\left[\frac{1}{24}\left(1-\frac{2}{9\sqrt{3}}\right)\right]\mathcal{H}^{\mu\nu\lambda}\mathcal{H}_{\mu}{}^{\rho\sigma}\mathcal{H}_{\nu\rho}{}^{\kappa}\mathcal{H}_{\lambda\sigma\kappa}+\kappa^{2}\left(-\frac{1}{24}+\frac{\sqrt{3}}{648}\right)\mathcal{H}^{\mu\lambda\rho}\mathcal{H}_{\nu\lambda\rho}\mathcal{H}^{\nu\kappa\sigma}\mathcal{H}_{\mu\kappa\sigma}$
$\phantom{\int}+\kappa^{2}\left(\frac{1}{144}\right)(\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda})^{2}\Bigg\}\sqrt{-g}\,d^{D}x.$ (39)
Comparison with (16) shows that no choice of $B_{3},B_{4}$ can yield (39). (A particularly easy way to see this is to try to match both $\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}$ and $\mathcal{H}^{\mu\lambda\rho}\mathcal{H}_{\nu\lambda\rho}\mathcal{H}^{\nu\kappa\sigma}\mathcal{H}_{\mu\kappa\sigma}$. This, then, leads to contradicting solutions for the value of $B_{3}$). In other words, in $D>4$ dimensions, we have to choose between a unitary action and interpreting the Kalb–Ramond field strength as torsion [30, 31]. However, this is not necessarily the case in $D=4$. Here, due to over-antisymmetrization, the following dimensionally dependent identities hold:
@@
One way to confirm these identities is to replace $\mathcal{H}$ with its dual in $D=4$, i.e.,
@@
We may use these identities to simplify the actions (16) and (32) (post-compactification to (3+1)-dimensions) and obtain the following [^7] It should be noted that since the field redefinitions are performed in the $D=10$-dimensional low-energy heterotic string effective action, we should set $D=10$ in Equation (16). Then, we may proceed with the compactification to $D=4$ dimensions, and only afterwards may we apply the dimensionally dependent identities to the (3+1)-dimensional action. Note that the processes of field redefinition and compactification do not “commute” in the sense that if we compactify first and then carry out the field redefinitions in $D=4$, we obtain an action with different coefficients. For more details, see Appendix (B).:
@@
and [^8] Interestingly, the invariant $\bar{G}^{+}_{6}$ vanishes in $4D$ for a totally antisymmetric torsion. The invariant $\bar{G}^{+}_{4}$, being the Gauss–Bonnet combination with torsion (Euler density in Einstein–Cartan spacetimes), reduces to the ordinary Gauss–Bonnet combination in the absence of torsion (Euler density in Einstein (general relativity) spacetimes).
$S_{\text{GCS}}^{D=4}$ $=\int\bigg\{\frac{1}{\kappa^{2}}(A_{4}+A_{5})(R_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}-4R_{\mu\nu}R^{\mu\nu}+R^{2})+$
$+\frac{8}{3}A_{5}\left(\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}-\frac{1}{3}\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}R\right)\bigg\}\sqrt{-g}\,d^{4}x.$ (45)
@@
$S$ $=\int\left[\frac{1}{2\kappa^{2}}\bar{R}+\frac{\alpha^{\prime}}{\kappa^{2}}\left(A_{4}\bar{G}^{+}_{4}+A_{5}\bar{G}^{+}_{5}\right)\right]\sqrt{-g}\,d^{4}x$
$=\int\Big[\frac{1}{2\kappa^{2}}\bar{R}+\frac{\alpha^{\prime}}{\kappa^{2}}\Big[A_{4}\left(\bar{R}_{\mu\nu\lambda\rho}\bar{R}^{\lambda\rho\mu\nu}-4\bar{R}_{\mu\nu}\bar{R}^{\nu\mu}+\bar{R}^{2}\right)+$
$+A_{5}\left(\bar{R}_{\mu\nu\lambda\rho}\bar{R}^{\mu\nu\lambda\rho}-4\bar{R}_{\mu\nu}\bar{R}^{\mu\nu}+\bar{R}^{2}\right)\Big]\Big]\sqrt{-g}\,d^{4}x$
$=\int\bigg[\frac{1}{2\kappa^{2}}\bar{R}+\frac{\alpha^{\prime}}{\kappa^{2}}\bigg[\frac{1}{16}\bar{R}^{2}+\frac{3}{224}\left(\frac{17}{3}-\frac{22}{9\sqrt{3}}\right)\left(\bar{R}_{\mu\nu\lambda\rho}\bar{R}^{\lambda\rho\mu\nu}-4\bar{R}_{\mu\nu}\bar{R}^{\nu\mu}\right)$
$-\frac{3}{224}\left(1-\frac{22}{9\sqrt{3}}\right)\left(\bar{R}_{\mu\nu\lambda\rho}\bar{R}^{\mu\nu\lambda\rho}-4\bar{R}_{\mu\nu}\bar{R}^{\mu\nu}\right)\bigg]\bigg]\sqrt{-g}\,d^{4}x\,.$ (51)
In $D=4$, the Gauss–Bonnet combination is a total derivative, which we may ignore. Thus, the full effective action reads
$S$ $=\int\Big[\frac{1}{2\kappa^{2}}R-\frac{1}{6}\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}$
$+\alpha^{\prime}\left(\frac{11}{126\sqrt{3}}-\frac{1}{28}\right)\left(\mathcal{H}_{\mu\lambda\rho}\mathcal{H}_{\nu}{}^{\lambda\rho}R^{\mu\nu}-\frac{1}{3}\mathcal{H}_{\mu\nu\lambda}\mathcal{H}^{\mu\nu\lambda}R\right)\Big]\sqrt{-g}\,d^{4}x.$ (52)
Since $\bar{G}^{+}_{4}$ is the Gauss–Bonnet combination in an Einstein–Cartan manifold, it is easy to see, from (45) and (III), that the effective string action is characterized by a single quadratic curvature invariant, namely $\bar{G}^{+}_{5}$, which provides the non-vanishing $\mathcal{H}\mathcal{H}R$ terms. In the next section, we shall discuss the emergence of dynamical axion fields from the totally antisymmetric torsion interpretation of the Green–Schwarz-modified Kalb–Ramond field strength (4), following [23] (or [2] in the StRVM context).
@@ IV Axion-Like Fields from the $\mathcal{H}$-Torsion
To this end, we first rewrite the $D=4$ effective action (III) in a more convenient form, with a generic coefficient of the higher-order corrections, as
@@
$A=\alpha^{\prime}\left(\frac{11}{126\sqrt{3}}-\frac{1}{28}\right)\simeq 0.015\,\alpha^{\prime}\,$, (54)
in the heterotic string case. The reader should recall, from the discussion in the previous section, that this value of the coefficient $A$ arose by taking the original action (10) of [23], with coefficients matching the string scattering amplitudes [28] and requiring a unitary action. Then, upon performing field redefinitions and reduction (compactification) to $D=4$, one arrives at (44). Finally, matching with the most general unitary action with a torsionful connection (45), one obtains (53).
In $D=4$, the Bianchi identity (7) can be contracted with the Levi–Civita tensor $\eta_{\mu\nu\lambda\rho}$, so that it can be expressed, in tensor notation, as the scalar identity
@@
The Levi–Civita tensor in $D=4$ is defined as [24]
@@
where $\epsilon$ denotes the Levi–Civita symbol (tensor density) with the conventions $\epsilon_{0123}=+1$,  $\epsilon^{0123}=-1,$ etc. Furthermore, the symbol (~)($\widetilde{\cdots}$) over the curvature and gauge field strength tensors denotes the corresponding duals, defined as
@@
We can now impose the Bianchi identity (55) as a constraint in the path integral of the action (53), following [23, 2]. We accomplish this by inserting the identity as a delta functional in the path integral:
@@
We can express this delta functional as a path integral of a pseudoscalar Lagrange multiplier field b(x)b(x) [^9] The partition function (58) is parity ($P$)-even, while the CS anomaly $\mathcal{G}(\boldsymbol{\omega},\boldsymbol{A})$ is parity-odd. Hence, the field $b(x)$ must be pseudoscalar so that the right-hand side of (59) below is parity-even, thus matching the even-parity nature of its left-hand side.:
@@
where we perform a partial integration to move the partial derivative to $b$ and discard the total derivative term. Thus, our effective action becomes
@@
This action is still quadratic in $\mathcal{H}$ and, therefore, can be integrated out analytically. To this end, we first write it in the following form:
@@
and we can proceed to integrate out $\mathcal{H}$. We perform a Wick rotation to a Euclidean version of the path integral as follows:
@@
So, $\det K=C\det Q$, where $C$ is a (formally infinite) field-independent “constant”, with no physical significance that can be absorbed in the path integral normalization constant $\mathcal{N}$, and thus canceled when one computes normalized correlation functions that are physically relevant.
@@
such that the overall factor of $\frac{1}{3}$ gets absorbed into the path integral normalization constant. Then, we can use the identity
@@
where the trace tr is over the internal (tensor) indices. In our case, the function $f$ is the natural logarithm and thus
@@
Thus, we calculate the (divergent) one-loop correction at order $O(A)\sim O(a^{\prime})$ to be
@@
We can then Wick rotate back to Minkowski space and write the full path integral as
@@
Finally, we shall determine the inverse of $K^{\mu\nu\lambda}{}_{\alpha\beta\gamma}$. Schematically, we have that
@@
Since $A\propto a^{\prime}$, we only keep the $\propto A$ term in the expansion. Therefore, we have the following order in $\alpha^{\prime}$:
@@
Before proceeding with calculating $JK^{-1}J$, and thus the resulting effective action in terms of metric and axion $b$ fields, we should now give, for completeness, the modification of (1) due to the higher-order corrections of the effective action (53) proportional to the coefficient $A$ ((54)). This can be achieved by studying the saddle points of the corresponding path integral over the $\mathcal{H}$-field for the actions (61) and (62). The latter corresponds to the classical equation of motion of the action with respect to the field strength $\mathcal{H}_{\mu\nu\lambda}$, which reads
@@
where $J_{\alpha\beta\gamma}$ is given in (62),  the inverse $(K^{-1})_{\mu\nu\lambda}^{\quad\,\,\,\,\alpha\beta\gamma}$ in (86) and (84), and the linear order in $A\propto\alpha^{\prime}$ (cf. (54)). After some straightforward algebraic manipulations, we finally arrive at
@@
Then, we can use the Schouten identity, namely the fact that over-antisymmetrized objects vanish, as follows:
@@
which is consistent with the generic result that a three-form in (3+1)-dimensional spacetimes is dual to a vector $\mathcal{V}^{\mu}$ in  the sense of the relation
@@
Indeed, from (IV) and (91), we can calculate the dual vector as
@@
For $A=0$, this yields (1), while for $A\neq 0$, we observe that one obtains modifications of the duality relation, which amount to a direction-dependent term proportional to components of the Ricci tensor contracted with $\partial^{\mu}b$. In the context of our StRVM, we are interested in (3+1)-dimensional Einstein spaces, for which $R_{\mu\nu}=\frac{R}{4}\,g_{\mu\nu}$,  which (90) reduces to
@@
This amounts to a modification of (1) by a global rescaling factor depending on the torsion-free scalar curvature of the maximally symmetric Einstein spacetime. As we shall discuss later on in this article, the quantity $|A\,R|\ll 1$ during the RVM inflationary scenario, so any singular behavior is avoided in this case.
@@
Therefore, we finally have
@@
The resulting effective action is therefore given by
@@
Upon rescaling $b\rightarrow\frac{1}{3\sqrt{2}}b$, we finally obtain an effective action with canonical kinetic terms for the KR axion field [^10] We follow here the conventions and normalizations of [23], which differ from those of [2] by a rescaling of $\alpha^{\prime}$ by a factor of $\sqrt{3}$.:
@@
On replacing $\delta^{(4)}(0)\sim\Lambda^{4}$, where $\Lambda$ is a UV cutoff energy scale, we observe that the coefficient of the Einstein–Hilbert (scalar curvature) term in the effective action (99) is positive, so unitarity is guaranteed.
In string effective field theories, the UV cutoff is the string mass scale
@@
whose value is bounded from above by $M_{\rm Pl}$ [19]. Using (100), the final form of the string-inspired effective action (99), containing terms of the fourth order in derivatives and respecting the unitarity and torsion interpretation of $H_{\mu\nu\rho}$, reads
@@
Therefore, we observe that the effects of keeping the complete four-derivative $\mathcal{O}(\alpha^{\prime})$ terms in the effective action of the bosonic massless string multiplet, under the assumption of constant dilaton, are the following:
@@
  • Introduction of a non-minimal derivative coupling of the KR axion b with the Ricci tensor, which maintains the shift symmetry of the axion, as it should, given that such terms originate from a path integration of the $H$-field strength, which is classically related to $b$ via (1).
It is worth noting that when expressed in terms of the dimensionless $\widetilde{b}(x)\equiv\kappa\,b(x)$ axion field, the coefficients of the four-spacetime derivative terms, that is, the CS anomalous terms and the non-minimal derivative coupling term, are both proportional to order $\alpha^{\prime}/\kappa^{2}\lesssim 1$. However, the relative magnitude of these terms can only be determined once we know the corresponding contributions at specific cosmological eras, which we shall carry out in the next section. Since we are working in an EFT/perturbative approach, the ghosts introduced by both the CS and the non-minimal derivative coupling term can be ignored as long as we stay below the UV cutoff [7, 8]. We now proceed to calculate the equations of motion. It will be convenient in what follows to use the following generic form of our effective action (101):
@@
For completeness we note that on using the Ricci-tensor identity